# Revisiting the Fundamental Planes of Black Hole Activity for Strong Jet Sources

### v2

@@ 2 Sample
To obtain the largest possible sample of strong jet source, we cross-referenced the FSRQs’ catalogue from (Hovatta et al., 2009), BL Lac objects’ catalog from Ye & Fan (2021) and Wu et al. (2014) with the synchrotron peak frequency $\log\nu_{\rm{peak}}$ of Ajello et al. (2022) and Xiong et al. (2015). Besides, we also add five LBLs that were not used in their works, their 5 GHz core radio flux densities ($F_{\rm{C,5GHz}}$) taken from NED [^1] https://ned.ipac.caltech.edu/ or Yuan & Wang (2012), their 5 GHz Doppler factors $\delta$ are estimated by using Equation (6) in Ye & Fan (2021). In total, we have collected $50$ FSRQs, $51$ LBLs, and $18$ IBLs for our strong jet sample; they have the available core radio flux density, X-ray flux density, BH mass, $\log\nu_{\rm{peak}}$ and Doppler factor $\delta$. It is noting that we focus only on the strong jet sources. Therefore, the radio-loudness $R$ of blazars are calculated to ensure that they are all RL-AGNs, where the boundary of the radio-loudness dichotomy of $\log R=1.23$ (ref Figure 6 in Zhang et al., 2021) is adopted in this work. In order to obtain the values of radio-loudness, we searched the optical flux densities at $\rm 4400\,\mbox{\AA}$ band in NED and CDS [^2] https://simbad.u-strasbg.fr/simbad/. For the sources without $F_{\rm 4400\,\mbox{\AA}}$, the other optical bands are extrapolated to $\rm 4400\,\mbox{\AA}$ assuming a spectral index of $\alpha_{o}=-0.5$ ($F\propto\nu^{\alpha}$, see Bariuan et al., 2022).
@@ 2.2 The 2–10 keV X-Ray Luminosity
In this paper, the 2–10 keV X-ray Luminosities are calculated by using 2-10 keV X-ray flux densities ($F_{\rm{2-10\,keV}}$). For our blazars, the X-ray observations are mainly detected by Swift, Chandra and XMM-Newton, their X-ray flux densities are taken from NED and the existing literatures. In order to minimize the observational errors, we preferred the X-ray data from Chandra/XMM-Newton if the source has the X-ray flux density that was detected by Chandra/XMM-Newton, owning to the Chandra/XMM-Newton have a higher spatial resolution. In addition, in order to avoid the errors caused by converting through another approximate wave band, we preferred the $F_{\rm{2-10\,keV}}$. For the sources without the available $F_{\rm{2-10\,keV}}$, the other approximate wavebands (e.g., 0.5–7 keV, 0.3–8 keV, 0.3–10 keV) are extrapolated to $F_{\rm{2-10\,keV}}$ by using the available power law photon index $\Gamma$ ($F_{\nu}\propto\nu^{1-\Gamma}$):

### v3

@@ 2 Sample
To obtain the largest possible sample of strong jet source, we cross-referenced the FSRQs’ catalogue from (Hovatta et al., 2009), BL Lac objects’ catalog from Ye & Fan (2021) and Wu et al. (2014) with the synchrotron peak frequency $\log\nu_{\rm{peak}}$ of Ajello et al. (2022) and Xiong et al. (2015). Besides, we also add five LBLs that were not used in their works, their 5 GHz core radio flux densities ($F_{\rm{C,5GHz}}$) taken from NASA/IPAC Extragalactic Database (NED) [^1] https://ned.ipac.caltech.edu/ or Yuan & Wang (2012), their 5 GHz Doppler factors $\delta$ are estimated by using Equation (6) in Ye & Fan (2021). In total, we have collected $50$ FSRQs, $51$ LBLs, and $18$ IBLs for our strong jet sample; they have the available core radio flux density, X-ray flux density, BH mass, $\log\nu_{\rm{peak}}$ and Doppler factor $\delta$. It is noting that we focus only on the strong jet sources. Therefore, the radio-loudness $R$ of blazars are calculated to ensure that they are all RL-AGNs, where the boundary of the radio-loudness dichotomy of $\log R=1.23$ (ref Figure 6 in Zhang et al., 2021) is adopted in this work. In order to obtain the values of radio-loudness, we searched the optical flux densities at $\rm 4400\,\mbox{\AA}$ band in NED and SIMBAD Astronomical Database - CDS [^2] https://simbad.u-strasbg.fr/simbad/. For the sources without $F_{\rm 4400\,\mbox{\AA}}$, the other optical bands are extrapolated to $\rm 4400\,\mbox{\AA}$ assuming a spectral index of $\alpha_{o}=-0.5$ ($F\propto\nu^{\alpha}$, see Bariuan et al., 2022).
@@ 2.2 The 2–10 keV X-Ray Luminosity
In this paper, the 2–10 keV X-ray Luminosities are calculated by using 2-10 keV X-ray flux densities ($F_{\rm{X,2-10}}$). For our blazars, the X-ray observations are mainly detected by Swift, Chandra and XMM-Newton, their X-ray flux densities are taken from NED and the existing literatures. In order to minimize the observational errors, we preferred the X-ray data from Chandra/XMM-Newton if the source has the X-ray flux density that was detected by Chandra/XMM-Newton, owning to the Chandra/XMM-Newton have a higher spatial resolution. In addition, in order to avoid the errors caused by converting through another approximate wave band, we preferred the $F_{\rm{X,2-10}}$. For the sources without the available $F_{\rm{X,2-10}}$, the other approximate wavebands (e.g., 0.5–7 keV, 0.3–8 keV, 0.3–10 keV) are extrapolated to $F_{\rm{2-10\,keV}}$ by using the available power law photon index $\Gamma$ ($F_{\nu}\propto\nu^{1-\Gamma}$):